# Explainable AI for computational pathology identifies model limitations and tissue biomarkers

### v2

@@ Methods
The degree to which tumor regions influence ABMIL models for metastasis detection remains unclear. To test the necessity of tumor regions, all of the tumor in the 49 tumor-positive specimens was removed, and change in model outputs was recorded. The embeddings of all patches intersecting with expert tumor annotations were removed. See Supplementary Fig. 13 for the histogram of the number of patches removed from each WSI. Once all of the tumor patches were removed from the bag of embeddings, the specimen was called “negative” for metastasis. The modified bags of embeddings were run through the model, and outputs were recorded. The true negative rate (specificity) was calculated as the ratio of true negative detections to all negative samples. In this case, as all samples were negative, the true negative rate was the proportion of specimens called negative by the model. This was done for all patch embedding tested in the present report.
@@
The extent to which tumor size affects specimen-level metastasis detection is incompletely understood. Conventionally, this analysis is limited to existing specimens. We explore a more rich set of tumor sizes using counterfactual examples. First, we evaluated the effect of a single tumor region in normal and metastatic specimens. The tumor region was taken from specimen test_001 at the coordinates (37878, 63530, 38444, 64096), indicating minimum X, minimum Y, maximum X, and maximum Y. For normal specimens, we added the embedding of this one patch into each of the 80 normal specimens and fed these bags of embeddings to the ABMIL model. Sensitivity was measured as the proportion of positive model predictions. We also evaluated this in the context of positive specimens. First, all tumor patches intersecting with expert tumor annotations were removed, and the single patch embedding was added to the bags of embeddings. 48 positive samples were used – the specimen that the patch came from was not included. Sensitivity was measured as the proportion of positive predictions.
@@
Last, the size of tumor was evaluated by sampling increasing numbers of tumor patches. First, all tumor patches intersecting with expert tumor annotations were removed. Then, tumor patches fully contained by the annotations were randomly sampled and added back to the bag of embeddings. This was evaluated over multiple numbers of sampled patches (i.e., 1, 2, 4, 8, 16, 32, 64). Sensitivity was evaluated as the proportion of positive predictions.
@@
Upon inspection of attention maps for the CTransPath-based metastasis detection model, adipose regions had high attention in a false negative, leading us to hypothesize that adipose regions were driving the false negative in that specimen. Attention alone could not allow us to address this hypothesis, but HIPPO could. The adipose regions were annotated in QuPath. Patches that intersected with the adipose region were removed, while ensuring that no tumor patches were removed. To measure the effect of this adipose tissue in other specimens, patches intersecting with the adipose annotation were added to the other 48 metastasis-positive slides, and the number of changes from true positive to false negative were recorded.
@@
We sought to evaluate how HIPPO can uncover shortcut learning and how it compares to attention in this regard. To do this, we modified the normal specimens in the CAMELYON16 dataset to include a blue square (hexadecimal color code #284283). This is meant to mimic a plausible real world scenario in which a pathologist marked certain slides with a blue pen. In practice, we embedded one blue square of using the UNI model [20] and replicated that embedding 36 times to create a blue region. The embeddings of this blue region were concatenated with the patch embeddings of normal specimens. The specimens with metastasis were not modified. We reasoned that the ABMIL model would learn to distinguish normal from metastatic specimens by the presence of a blue region. To assess whether tumor regions were affecting model predictions in positive specimens, we removed all patches intersecting with tumor annotations in positive specimens and recorded model outputs. To visualize attention maps, we saved patch-wise attention weights in GeoJSON format and visualized the maps in QuPath [94]. We also used the search strategy HIPPO-search-high-effect to identify the regions with highest effect sizes de novo. We also did this using a UNI-based ABMIL model trained on the original, unaltered CAMELYON16 dataset, trained using the same hyperparameters and random seed.
@@
We sought to compare the effectiveness of attention and HIPPO for identifying tissue regions related to predicted prognosis. TCGA BRCA and SKCM data were used in these experiments. For attention, regions assigned the top 1% of attention scores were selected. For HIPPO, the search strategy HIPPO-search-high-effect was used to identify the regions most contributing to high risk in high-risk specimens, and the search strategy HIPPO-search-low-effect was used to identify the regions most contributing to low risk in low-risk specimens. Low and high risk were defined as the first and fourth quartiles of predicted risk scores, respectively. The first 1% of patches identified by the HIPPO search algorithms were selected for evaluation. To quantify the effect of the selected regions on predicted prognosis, we calculated the difference between the predicted prognosis on the original specimens and the predicted prognosis on the specimens with the selected regions removed.
Positive values indicated that the regions contributed to higher risk, and negative values indicated that the regions contributed to lower risk. Independent t-tests were used to assess significance of differences between attention and HIPPO.
@@
In prognostic models, we measured the effects of tumor-infiltrating lymphocytes (TILs) on model behavior. The number of TILs was quantified using the same approach as Ref. [50]. Briefly, HoVer-Net [96] was used to outline and label the nuclei in TCGA BRCA and SKCM WSIs. The model labels nuclei as one of six categories: tumor epithelium, lymphocyte, stroma, necrosis, normal epithelium, and unknown. Each was called TIL-positive if it contained more than 20 cells, more than 10 immune cells, and more than 5 tumor cells. In TCGA BRCA, HoVer-Net failed for 12 WSIs, some of which were missing pixel spacing information.
We measured the effect of TIL patches on predicted prognosis in TCGA BRCA AND SKCM by either removing TILs from low-risk specimens or adding TILs to high-risk specimens, where low-risk was defined as samples in the first quartile of predicted risk and high-risk were samples in the fourth quartile of predicted risk. The predicted prognoses were compared before and after the intervention. To evaluate the sufficiency of TILs for predicting low risk, we added TIL patches from low-risk specimens to high-risk specimens. Risk predictions of the model were recorded, and differences were tested using paired t-tests. To assess the necessity of TIL regions, we removed TIL-positive patches from low risk specimens and measured risk predictions. Differences were tested using paired t-tests.
@@
Autologous TIL therapy is a promising immunotherapy. We explored how HIPPO could be used for hypothesis generation in the context of autologous TILs in high-risk SKCM specimens (n=67). We sought to assess the degree to which prognostic ABMIL models are effected by the number of TILs in a specimen. We do not claim to assess the efficacy of autologous TILs through HIPPO. The embeddings of TIL-positive regions were replicated , , , or , and the change in predicted risk was measured:
Negative values indicated that the addition of TILs decreased risk. The change in risk from baseline was assessed using paired t-tests.
@@
We analyzed histopathology specimens with known IDH mutations but misclassified as negative by the models (i.e., false negative, FN). The objective was to evaluate whether removing the top of patches, identified by attention or HIPPO search, could increase the likelihood of a correct classification. To reduce running time of experiments, we analyzed FN specimens with fewer than patches (n=38). For each specimen, the softmax probability for IDH mutation was recorded. Next, the of patches with the highest attention scores were removed, and the updated model probability was documented. In addition, the top of patches identified by HIPPO-search-high-effect were removed, and the updated model probability was recorded (search algorithm described below). The final IDH mutation probabilities after the different methods of patch removal were compared across baseline, attention, and HIPPO conditions. Cohen’s d was used to assess effect sizes, and independent t-tests evaluated statistical significance between conditions. The experiment was conducted exclusively on specimens from the test set of each fold during the 10-fold Monte Carlo cross-validation, ensuring that no specimens from the training or validation sets were included.
The HIPPO-search-high-effect algorithm iteratively removed individual patches, replacing each patch after recording its impact on mutation probability. The patch whose removal most increased the probability of IDH mutation was permanently excluded, with this process repeated until of patches were removed. The search was performed using NVIDIA A100 GPUs.
@@
We evaluated the extent to which attention and HIPPO search identified patches driving positive predictions, focusing on specimens with positive predictions in the first Monte Carlo cross-validation fold (n=25). The HIPPO-search-high-effect algorithm was applied, removing patches in batches of 10 per iteration to accelerate runtime. Patches were ranked by their impact on IDH mutation probabilities, with the 10 patches whose removal most reduced the prediction probability excluded at each step. The IDH mutation probability was recorded after each batch, continuing until of patches were removed. A similar approach was used for attention, removing patches in descending order of attention scores in batches of 10 and recording model outputs after each removal. Model outputs for each specimen were visualized as line plots, comparing results from HIPPO and attention. Heatmaps of attention and HIPPO outputs were viewed using QuPath for qualitative assessment.
To assess the efficiency of HIPPO in identifying critical patches, we conducted an experiment to determine how many patches needed to be removed to reduce the model probability below 0.4, hypothesizing that HIPPO would require fewer patches than attention due to its ability to better capture the most important regions driving predictions. The HIPPO and attention outputs from above were used in this experiment. The ratio of patches required (i.e., number of patches required normalized by number of patches in specimen) were compared between attention and HIPPO using independent t-test and visualized using a scatter plot.
@@
A Python package implementing HIPPO is available at https://github.com/kaczmarj/HIPPO and is licensed under the terms of the 3-Clause BSD License. HIPPO documentation is published at https://github.com/kaczmarj/HIPPO under the terms of the Creative Commons Attribution-NonCommercial-ShareAlike 4.0 International copyright license (CC BY-NC-SA 4.0). Model weights and inference code are available at the following repositories: UNI (https://huggingface.co/MahmoodLab/UNI), REMEDIS (https://github.com/google-research/medical-ai-research-foundations), Phikon (https://huggingface.co/owkin/phikon), CTransPath (https://github.com/Xiyue-Wang/TransPath), and RetCCL (https://github.com/Xiyue-Wang/RetCCL). Model weights for the models trained for this report will be deposited to online repositories.
Acknowledgements
This research was supported by National Science Foundation (NSF) grant IIS2212046, National Institutes of Health (NIH) grant UH3CA225012, and Stony Brook Profund 2022 seed funding. JRK was also supported by the Medical Scientist Training Program at Stony Brook University and NIH grant T32GM008444 (NIGMS). We would also like to acknowledge the Department of Biomedical Informatics at Stony Brook University and the Simons Center for Quantitative Biology at Cold Spring Harbor Laboratory.
@@ Acknowledgements
JRK and PKK conceived of the method and planned the experiments. JRK wrote the code and ran all experiments. JRK, PKK, and JHS interpreted the results of experiments. JHS and PKK supervised the project. JRK wrote the initial draft of the manuscript. All authors provided feedback on the manuscript and contributed to the final manuscript.

### v3

@@ Methods
The degree to which tumor regions influence ABMIL models for metastasis detection remains unclear. To test the necessity of tumor regions, all of the tumor in the 49 tumor-positive specimens was removed, and changes in model outputs were recorded. The embeddings of all patches intersecting with expert tumor annotations were removed. See Supplementary Fig. LABEL:fig:supp_hist_patches_with_mets for the histogram of the number of patches removed from each WSI. Once all of the tumor patches were removed from the bag of embeddings, the specimen was called “negative” for metastasis. The modified bags of embeddings were run through the model, and outputs were recorded. The true negative rate (specificity) was calculated as the ratio of true negative detections to all negative samples. In this case, as all samples were negative, the true negative rate was the proportion of specimens called negative by the model. This was done for all patch embedding tested in the present report.
@@
The extent to which tumor size affects specimen-level metastasis detection is incompletely understood. Conventionally, this analysis is limited to existing specimens. We explore a richer set of tumor sizes using counterfactual examples. First, we evaluated the effect of a single tumor region in normal and metastatic specimens. The tumor region was taken from specimen test_001 at the coordinates (37878, 63530, 38444, 64096), indicating minimum X, minimum Y, maximum X, and maximum Y. For normal specimens, we added the embedding of this one patch into each of the 80 normal specimens and fed these bags of embeddings to the ABMIL model. Sensitivity was measured as the proportion of positive model predictions. We also evaluated this in the context of positive specimens. First, all tumor patches intersecting with expert tumor annotations were removed, and the single patch embedding was added to the bags of embeddings. 48 positive samples were used – the specimen that the patch came from was not included. Sensitivity was measured as the proportion of positive predictions.
@@
Lastly, the size of tumor was evaluated by sampling increasing numbers of tumor patches. First, all tumor patches intersecting with expert tumor annotations were removed. Then, tumor patches fully contained by the annotations were randomly sampled and added back to the bag of embeddings. This was evaluated over multiple numbers of sampled patches (i.e., 1, 2, 4, 8, 16, 32, 64). Sensitivity was evaluated as the proportion of positive predictions.
@@
Upon inspection of attention maps for the CTransPath-based metastasis detection model, adipose regions had high attention in a false negative, leading us to hypothesize that adipose regions were driving the false negative in that specimen. Attention alone could not allow us to address this hypothesis, but HIPPO could. The adipose regions were annotated in QuPath and verified by two board-certified histopathologists . Patches that intersected with the adipose region were removed, while ensuring that no tumor patches were removed. To measure the effect of this adipose tissue in other specimens, patches intersecting with the adipose annotation were added to the other 48 metastasis-positive slides, and the number of changes from true positive to false negative were recorded.
@@
We sought to evaluate how HIPPO can uncover shortcut learning and how it compares to attention in this regard. To do this, we modified the normal specimens in the CAMELYON16 dataset to include a blue square (hexadecimal color code #284283). This is meant to mimic a plausible real-world scenario in which a pathologist marked certain slides with a blue pen. In practice, we embedded one blue square of using the UNI model [chen2024uni] and replicated that embedding 36 times to create a blue region. The embeddings of this blue region were concatenated with the patch embeddings of normal specimens. The specimens with metastasis were not modified. We reasoned that the ABMIL model would learn to distinguish normal from metastatic specimens by the presence of a blue region. To assess whether tumor regions were affecting model predictions in positive specimens, we removed all patches intersecting with tumor annotations in positive specimens and recorded model outputs. To visualize attention maps, we saved patch-wise attention weights in GeoJSON format and visualized the maps in QuPath [bankhead2017qupath]. We also used the search strategy HIPPO-search-positive-effect to identify the regions with the highest effect sizes de novo. We also did this using a UNI-based ABMIL model trained on the original, unaltered CAMELYON16 dataset, trained using the same hyperparameters and random seed.
@@
We sought to compare the effectiveness of attention and HIPPO for identifying tissue regions related to predicted prognosis. TCGA BRCA and SKCM data were used in these experiments. For attention, regions assigned the top 1% of attention scores were selected. For HIPPO, the search strategy HIPPO-search-positive-effect was used to identify the regions most contributing to high risk in high-risk specimens, and the search strategy HIPPO-search-negative-effect was used to identify the regions most contributing to low risk in low-risk specimens. Low and high risk were defined as the first and fourth quartiles of predicted risk scores, respectively. The first 1% of patches identified by the HIPPO search algorithms were selected for evaluation. To quantify the effect of the selected regions on predicted prognosis, we calculated the difference between the predicted prognosis on the original specimens and the predicted prognosis on the specimens with the selected regions removed.
Positive values indicated that the regions contributed to higher risk, and negative values indicated that the regions contributed to lower risk. Independent two-sided t-tests were used to assess the significance of differences between attention and HIPPO.
@@
In prognostic models, we measured the effects of tumor-infiltrating lymphocytes (TILs) on model behavior. The number of TILs was quantified using the same approach as Ref. [chen2022porpoise]. Briefly, HoVer-Net [graham2019hovernet] was used to outline and label the nuclei in TCGA BRCA and SKCM WSIs. The model labels nuclei as one of six categories: tumor epithelium, lymphocyte, stroma, necrosis, normal epithelium, and unknown. Each was called TIL-positive if it contained more than 20 cells, more than 10 immune cells, and more than 5 tumor cells. In TCGA BRCA, HoVer-Net failed for 12 WSIs, some of which were missing pixel spacing information.
We measured the effect of TIL patches on predicted prognosis in TCGA BRCA and SKCM by either removing TILs from low-risk specimens or adding TILs to high-risk specimens, where low-risk was defined as samples in the first quartile of predicted risk and high-risk were samples in the fourth quartile of predicted risk. The predicted prognoses were compared before and after the intervention. To evaluate the sufficiency of TILs for predicting low risk, we added TIL patches from low-risk specimens to high-risk specimens. Risk predictions of the model were recorded, and differences were tested using paired two-sided t-tests. To assess the necessity of TIL regions, we removed TIL-positive patches from low-risk specimens and measured risk predictions. Differences were tested using paired two-sided t-tests.
@@
Autologous TIL therapy is a promising immunotherapy. We explored how HIPPO could be used for hypothesis generation in the context of autologous TILs in high-risk SKCM specimens (n=67). We sought to assess the degree to which prognostic ABMIL models are affected by the number of TILs in a specimen. We do not claim to assess the efficacy of autologous TILs through HIPPO. The embeddings of TIL-positive regions were replicated , , , or , and the change in predicted risk was measured:
Negative values indicated that the addition of TILs decreased risk. The change in risk from baseline was assessed using paired two-sided t-tests.
@@
We analyzed histopathology specimens with known IDH mutations but misclassified as negative by the models (i.e., false negative, FN). The objective was to evaluate whether removing the top of patches, identified by attention or HIPPO search, could increase the likelihood of a correct classification. To reduce running time of experiments, we analyzed FN specimens with fewer than patches (n=38). For each specimen, the softmax probability for IDH mutation was recorded. Next, the of patches with the highest attention scores were removed, and the updated model probability was documented. In addition, the top of patches identified by HIPPO-search-negative-effect were removed, and the updated model probability was recorded (search algorithm described below). The final IDH mutation probabilities after the different methods of patch removal were compared across baseline, attention, and HIPPO conditions. Cohen’s d was used to assess effect sizes, and independent two-sided t-tests evaluated statistical significance between conditions. The experiment was conducted exclusively on specimens from the test set of each fold during the 10-fold Monte Carlo cross-validation, ensuring that no specimens from the training or validation sets were included.
The HIPPO-search-positive-effect algorithm iteratively removed individual patches, replacing each patch after recording its impact on mutation probability. The patch whose removal most increased the probability of IDH mutation was permanently excluded, with this process repeated until of patches were removed. The search was performed using NVIDIA A100 GPUs.
For applying HIPPO to transformers, we used the Vision Transformer architecture which was adapted to the permutation-invariant MIL setting by excluding the positional embeddings [el2025stamp]. To extract attention scores from the ViT model, we considered two approaches: (1) Attention (Last), which uses the raw attention weights from the CLS token to each patch token in the final self-attention layer, (2) Attention (Rollout), which is computed by recursively multiplying CLS-centric attention across all the transformer layers, following prior work[abnar2020quantifying].
@@
We evaluated the extent to which attention and HIPPO search identified patches driving positive predictions, focusing on specimens with positive predictions in the first Monte Carlo cross-validation fold (n=25). The HIPPO-search-positive-effect algorithm was applied, removing patches in batches of 10 per iteration to accelerate runtime. Patches were ranked by their impact on IDH mutation probabilities, with the 10 patches whose removal most reduced the prediction probability excluded at each step. The IDH mutation probability was recorded after each batch, continuing until of patches were removed. A similar approach was used for attention, removing patches in descending order of attention scores in batches of 10 and recording model outputs after each removal. Model outputs for each specimen were visualized as line plots, comparing results from HIPPO and attention. Heatmaps of attention and HIPPO outputs were viewed using QuPath for qualitative assessment.
To assess the efficiency of HIPPO in identifying critical patches, we conducted an experiment to determine how many patches needed to be removed to reduce the model probability below 0.4, hypothesizing that HIPPO would require fewer patches than attention due to its ability to better capture the most important regions driving predictions. The HIPPO and attention outputs from above were used in this experiment. The ratio of patches required (i.e., number of patches required normalized by number of patches in specimen) were compared between attention and HIPPO using an independent two-sided t-test and visualized using a scatter plot.
@@
A Python package implementing HIPPO is available at https://github.com/kaczmarj/HIPPO and is licensed under the terms of the 3-Clause BSD License. HIPPO documentation is published at https://github.com/kaczmarj/HIPPO under the terms of the Creative Commons Attribution-NonCommercial-ShareAlike 4.0 International copyright license (CC BY-NC-SA 4.0). Model weights and inference code are available at the following repositories: UNI (https://huggingface.co/MahmoodLab/UNI), REMEDIS (https://github.com/google-research/medical-ai-research-foundations), Phikon (https://huggingface.co/owkin/phikon), CTransPath (https://github.com/Xiyue-Wang/TransPath), RetCCL (https://github.com/Xiyue-Wang/RetCCL), and Virchow2 (https://huggingface.co/paige-ai/Virchow2). The ViT model used in the IDH mutation classification experiment is available at the following repository: https://github.com/KatherLab/STAMP. Model weights for the models trained for this report will be deposited to online repositories.
Acknowledgments
This research was supported by National Science Foundation (NSF) grant IIS2212046, National Institutes of Health (NIH) grant UH3CA225012, R01EB035934, and Stony Brook Profund 2022 seed funding. JRK was also supported by the Medical Scientist Training Program at Stony Brook University and NIH grant T32GM008444 (NIGMS). This work was performed with assistance from CSHL Shared Resources, including the CSHL Animal & Tissue Imaging Shared Resource, which are supported by the Cancer Center Support Grant 5P30CA045508. We would also like to acknowledge the Department of Biomedical Informatics at Stony Brook University and the Simons Center for Quantitative Biology at Cold Spring Harbor Laboratory.
@@ Acknowledgments
JRK and PKK conceived of the method. JRK, CK, SG, SL, and PKK planned the experiments. JRK, CK, and SG wrote the code and ran all experiments. All authors interpreted the results of experiments. ZZ and DS provided expert annotations for select histopathology patches. SL, JHS, and PKK supervised the project. JRK, CK, SG, SL, JHS, and PKK contributed to the writing of the manuscript. All authors provided feedback on the manuscript and contributed to the final manuscript.